he Morris Corporation has $350,000 of debt outstanding, and it pays an interest rate of 9% annually. Morris's annual sales are $1.75 million, its average tax rate is 35%, and its net profit margin on sales is 5%. If the company does not maintain a TIE ratio of at least 6 to 1, then its bank will refuse to renew the loan, and bankruptcy will result. What is Morris's TIE ratio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Net profit margin=Net profit/Sales

Net profit=(1,750,000*5%)=$87500

EBIT(balance)(134,615.38+31500) 166,115.38(Approx).
Less:interest(350,000*9%) 31500
EBT(100%)(87500/0.65) 134,615.38
Less:tax@35%(134,615.38*35%) 47115.38
Net profit(65%) 87500

TIE ratio=EBIT/interest expense

=166,115.38/31500

=5.27(Approx).